[figure: 9.7 cm]

MAGNETISM AND ELECTRONIC STRUCTURE OF LaCoO_{3}

Z. Ropka^{a}, R. Michalski^{a,b} and R.J. Radwański^{a,b} aCenter for Solid State Physics, św. Filip 5, 31-150 Kraków bInst. of Physics, Pedagogical University, 30-084 Kraków, Poland.

LaCoO_{3} exhibits unusual magnetic properties intriguing the magnetic society by more than 30 years, after pioneering works of Goodenough (Phys.Rev. 155 (1967) 932). LaCoO_{3} show a diamagnetic low-temperature state that seems to transform with temperature into a magnetic state. It is often discussed in terms of low-spin and high-spin states. Recently an intemediate-spin state has been additionally postulated to exist (Phys.Rev. B 54 (1996) 5309). We have performed calculations of the fine electronic structure of the Co^{3+} ion in the perovskite-like structure of LaCoO_{3}. We have found that the diamagnetic state of LaCoO_{3} is related with the non-magnetic state of the Co^{3+} ion in the atomic scale. Moreover, we have found microscopic picture for the existence of the low-, intermediateand high-spin states of the cobalt ions and their transformation with temperature. Very strong correlation between the local magnetism, the local crystallographic symmetry (Jahn-Teller effect) and the fine electronic structure has been found. The present results confirm the importance of the spin-orbit interactions on the 3d-ion moment and can enlighten mechanisms responsible for the formation of 3d magnetic moments.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Z. Ropka

Address for correspondence : Center for Solid State Physics, św. Filip 5, 31-150 Kraków

Email address : sfradwan@cyf-kr.edu.pl